Review I Constants Suppose you take and hold a deep breath on a chilly day inhaling 1.7 L of air at 0° C. Assume that air pressure is 1.0 atm
How much heat must your body supply to warm the air to your internal body temperature of 37°C? Assume that for air C , consists mostly of nitrogen and oxygen. 29 1 J (K-mo as it Express your answer to two significant figures and include the appropriate units. Value Units
How much does the volume of the air increase as it is warmed? Express your answer to two significant figures and include the appropriate units. 砡 LA aue Units
Answer #1

By ideal gas equation,

P*V = n*R*T

So, moles of gas(n) = P*V/(R*T)

given, P = 1 atm

V = 1.7 L

T = 0 deg C = 273 K

R = 0.0821

So, n = 1*1.7/(0.0821*273)

n = 0.076

Since, heat transfer(Q) = n*C*dT

given,

C = 29.1 J/(K*mole)

dT = 37 - 0 = 37

So,  Q = 0.076*29.1*37

Q = 82 J

Part B.

Since pressure is constant,

So, by ideal gas equation,

Vi/Vf = Ti/Tf

here initially,

Ti = 273 K   &    Vi = 1.7 L

Tf = 273+37 = 310 K

Vf = ??

So,

Vf = (1.7*310)/273

Vf = 1.9 L

Change in Volume = Vf - Vi = 1.9 - 1.7 = 0.2 L

Please upvote.